OpenStudy (raiders88):

whAat is the y-intercept of the graph of the line given by the equation 3y+4x=15

OpenStudy (anonymous):

SO this equation given is in standard form, we need to put it in slope intercept, do you know what that is?

OpenStudy (anonymous):

the slope intercept form is: y = mx + b where m is slope and b is your y intercept. Here are the steps to get to that. 3y+4x = 15 We need to isolate x: -4x on both sides 3y = -4x + 15 /3 on both sides y = (-4/3)x + 5 So your y - intercept is 5

OpenStudy (raiders88):

thx i got a few more if u cn help

OpenStudy (anonymous):

actually just go ahead and post them here

OpenStudy (michele_laino):

please set x=0, and solve for y

OpenStudy (raiders88):

solve the system of equations by substitutions : x=-4y x+5y

hero (hero):

@twistnflip The proper way to help users on OS is to involve them in the problem solving process. Find out what they know and what they don't know first, then guide them through the problem solving process. User participation is a huge part of how OS should operate. Explaining the answer by posting all the steps is equivalent to giving the answer.

OpenStudy (anonymous):

I spoke to him in chatting and then just related what we where talking about here. Sorry

hero (hero):

The intention behind the question threads were to create an environment that encouraged a engaging conversation between the asker and the helper in a way that demonstrates the asker learning and comprehending concepts explained by the helper. The helper should ask the helper questions and give helpful clues and hints on how to solve the problem rather than answer the problem completely. If the asker isn't involved in the process then he or she did not learn anything. Active participation is the key.

OpenStudy (anonymous):

Got it, I will try to do better next time.

OpenStudy (raiders88):

megan..........................why

hero (hero):

We're trying to discourage helpers from just giving answers such as "The answer is A" or just posting all the steps. What we want to see is engaging conversation, askers being guided to the answer by actively participating in the problem solving process. We want helpers to identify where the asker is stuck and then give them helpful clues on how to proceed to the next step in the problem solving process. We want to see actual learning taking place.

OpenStudy (raiders88):

can u excuse me i got some personal businees to attend